OpenStudy (lilah):

When you multiply a function by -1, what is the effect on its graph? A. The graph flips over the x-axis B. The graph flips over the y-axis C. The graph flips over the line y=x

OpenStudy (lilah):

@satellite73

satellite73 (satellite73):

what as up is not down, what was down is now up

satellite73 (satellite73):

for example, look at \(y=x^2\) and \(y=-x^2\)

OpenStudy (lilah):

A??

satellite73 (satellite73):

yes

OpenStudy (lilah):

Thanks :)

satellite73 (satellite73):

yw